OpenStudy (anonymous):

How do you find the explicit definition for the sequence -2,1,4,7?

jimthompson5910 (jim_thompson5910):

Notice how you're adding 3 to each term to get the next term (eg: -2+3 = 1 or 4+3 = 7) So the common difference is 3

jimthompson5910 (jim_thompson5910):

The first term is -2

OpenStudy (anonymous):

I know that, I'm just not sure how to set it up

jimthompson5910 (jim_thompson5910):

First term: a1 = -2 Common difference: 3 Plug these into an = a1 + d(n-1) to get the explicit rule an = a1 + d(n-1) an = -2 + 3(n-1) an = -2 + 3n-3 an = 3n-5

OpenStudy (anonymous):

So that's the formula you plug it into?

jimthompson5910 (jim_thompson5910):

So let's say we wanted to find the 3rd term. That means n = 3 an = 3n - 5 a3 = 3*3 - 5 a3 = 9 - 5 a3 = 4 and 4 is indeed the 3rd term. I recommend you check the others.

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

Wait Im lost, where did the 5 come from?

jimthompson5910 (jim_thompson5910):

-2 - 3 is -5

OpenStudy (anonymous):

Oh okay

jimthompson5910 (jim_thompson5910):

look above in my steps and you'll see me jump from an = -2 + 3n-3 to an = 3n-5

OpenStudy (anonymous):

So would the definition be an = -2 + (n+3) ?

jimthompson5910 (jim_thompson5910):

no the definition is the last line an = 3n - 5

jimthompson5910 (jim_thompson5910):

basically the nth term is 3n - 5

jimthompson5910 (jim_thompson5910):

where n is some positive whole number

jimthompson5910 (jim_thompson5910):

Usually you have to simplify and combine like terms as much as possible

OpenStudy (anonymous):

Im still kind of confused because these are the choices I have

jimthompson5910 (jim_thompson5910):

oh they didn't simplify

jimthompson5910 (jim_thompson5910):

how odd

jimthompson5910 (jim_thompson5910):

then it would be an = -2 + 3(n-1)

jimthompson5910 (jim_thompson5910):

all I did was plug in a1 = -2, d = 3 into an = a1 + d(n-1)

OpenStudy (anonymous):

ohhhhhhhh ok

OpenStudy (anonymous):

Ok i understand more now

jimthompson5910 (jim_thompson5910):

again, the first term is a1 = -2, the common difference is d = 3

jimthompson5910 (jim_thompson5910):

that's good